I am 6 years older than my sister.How old will I be when I am 3 times older than she?
Answer by nerdybill(7384)

I am 6 years older than my sister.How old will I be when I am 3 times older than she?
.
Let x = sister's age
then
x-6 = "my age"
.
3(x-6) = x
3x-18 = x
2x-18 = 0
2x = 18
x = 9 yrs (sister's age)
.
"my age":
x-6 = 9-6 = 3 years old (solution)